### 31 Jan 2024 v1

@@ Abstract
Context:AD Leo is a young and active M dwarf with high flaring rates across the X-ray to radio bands. Flares accelerate particles in the outer coronal layers and often impact exo-space weather. Wide-band radio dynamic spectra let us explore the evolution of particle acceleration activity across the corona. Identifying the emission features and modelling the mechanisms can provide insights into the possible physical scenarios driving the particle acceleration processes.
@@
Methods:A python based module, named VISAD, was developed to obtain the visibility averaged wide-band dynamic spectra. Direct imaging was also performed with different frequency-time averaging. Based on existing observational results on AD Leo and on solar active region models, radial profiles of electron density and magnetic fields were derived. Applying these models, we explore the possible emission mechanisms and magnetic field profile of the flaring active region.
Results:The star displayed high brightness temperature ( 10 – 10 K) throughout the observation. The emission was also nearly 100% left circularly polarised during bursts. Post flare phase is characterised by a highly polarised (60 - 80%) solar-like type IV burst confined above 700 MHz.
Conclusions:The flare emission favors a Z -mode or a higher harmonic X-mode electron cyclotron maser emission mechanism. The 700 MHz post-flare activity is consistent with a type-IV radio burst from flare-accelerated particles trapped in magnetic loops, which could be a coronal mass ejection (CME) signature. This is the first solar-like type-IV burst reported on a young active M dwarf belonging to a different age - related activity population (‘C’ branch) compared to the Sun (‘I’ branch). We also find that, a multipole expansion model of the active region magnetic field better accounts for the observed radio emission than a solar-like active region profile.
@@
Low-mass main-sequence stars of spectral type M (M dwarfs) make up the most numerous stellar type in our galaxy. They also have the highest occurrence rate of Earth-like exoplanets in their habitable zones (Bashi et al., 2020). Meanwhile, M-dwarfs form the most active population among cool main-sequence stars with relatively high values of flaring rates, X-ray to bolometric flux ratio, and other activity metrics (see Donati & Landstreet, 2009, for a review). Moreover, the younger stars are more active with stronger magnetic fields and faster rotation rates (e.g., Vidotto et al., 2014; Lund et al., 2020). Stellar flares cause energetic space-weather events including the release of energetic particles, intense high-energy radiation and large-scale eruptions/outflows, which threaten the stability of planetary atmospheres and hence the habitability. So, it is important to understand the mechanism, evolution and space-weather consequences of M dwarf flares.
@@
Flare emission in the metrewave band is sensitive to various coherent emission signatures from flare-accelerated electrons either trapped in low-coronal magnetic field structures or streaming across coronal layers (e.g., Wild, 1970; McLean & Labrum, 1985; Bastian, 1990; Melrose, 2009, 2017a; Zic et al., 2020; Vedantham, 2021). The two prominant coherent emission mechanisms discussed in the literature include the electron cyclotron maser (ECME) emission and the plasma emission (see, Dulk, 1985; Güdel, 2002; Melrose, 2017a, for an overview). ECME dominates when the ratio of plasma frequency to gyro-frequency (/) at the flare source region is less than 1.5 (Melrose & Dulk, 1982; Melrose et al., 1984a; Dulk, 1985; Dulk & Winglee, 1987). When /3, the favoured emission mechanism is plasma emission (e.g., Ginzburg & Zhelezniakov, 1958; Tsytovich & Kaplan, 1969; Melrose & Sy, 1972; Cairns & Melrose, 1985; Melrose, 2009). Depending on the emission mechanism, the coherent emission frequencies correspond to either the fundamental mode or the harmonics of the local plasma frequency () or gyro-frequency (), letting us infer either the local density or magnetic field strength, respectively. Apart from the flare emission, M dwarfs are also known to display non-flaring (or non-impulsive) continuum emission with high brightness temperature () (e.g., White et al., 1989; Bastian, 1990). In a few GHz (¡2 GHz) bands, the continuum ranges from 10 - 10 K (Osten et al., 2008; Villadsen & Hallinan, 2019). In metrewavebands can be even higher (e.g., Bastian, 1994; Güdel, 2002; Villadsen & Hallinan, 2019). The continuum emission can form either from a gyrosynchrotron mechanism, or a combination of various coherent emission processes operating at multiple active regions (Bastian, 1996; Benz, 1996; Zaitsev et al., 1997).
@@
There have been many studies that explored radio emission from AD Leo, reporting several flaring events. While instrumental limitations forced earlier studies to use only a few disjoint narrow frequency bands (e.g., Spangler et al., 1974; Bastian et al., 1990), recent observations with continuous wide-band instruments has enabled spectroscopic explorations revealing rich variability in the dynamic spectrum (e.g., Osten et al., 2008; Villadsen & Hallinan, 2019; Zhang et al., 2023). These studies have revealed several features like fast pulsations at scales of ms - min and high drift rates in the dynamic spectrum, often related to the propagation of energetic particles or spatio-temporal evolution of plasma instabilities in the associated active regions (e.g., Bastian et al., 1990; Osten & Bastian, 2006; Osten et al., 2008; Zhang et al., 2023). However, much of these observations had been close to or above 1 GHz. Here we present a 8 h long monitoring of AD Leo across a wide sub-GHz band, which is expected to probe higher atmospheric layers than those probed by GHz bands. Though, the original study was planned with simultaneous observation spanning from X-ray to radio band, only the radio observations were successful due to various technical issues. We present the results from meterwave observations here.
@@ 2 Data and Methodology
AD Leo was observed with uGMRT (Gupta et al., 2017) in Band 4 (550 - 850 MHz) from 3 December 2021, 21:17:41 UT to 4 December 2021, 04:17:41 UT at 5 s and 100 kHz time and frequency resolution, respectively. The data were gathered in RR and LL polarisations, letting us explore the spectro-temporal evolution of the circular polarisation along with brightness temperature () or flux density. The data analysis was done using the functionalities in the interferometric analysis package, Common Astronomy Software Applications (CASA; CASA Team et al., 2022). The data were first flagged and calibrated using standard procedures. The flux calibrator was 3C147, and the phase calibrator was 1021+219. Portions of data affected by radio frequency interference from terrestrial sources and systematic instrumental effects were identified and flagged manually. The automatic flagging routine in CASA, namely tfcrop, was also used. This was followed by bandpass calibration using the flux calibrator, and then by time dependent gain calibration using the phase calibrator. The bandpass solution and the time dependent gain solution was then extrapolated to the target star field. The calibrated target field data was further flagged using the Rflag routine in CASA. Initial imaging was done using the CASA task tclean in mtmfs mode, with nterms set to 2 averaging over the entire observing time. This produced a mean model image for the entire spectral band. This was taken as the initial model and self-calibration was performed in an iterative manner (Cornwell & Fomalont, 1999) to improve this model. After the second iteration, an optimal model was obtained. So we stopped at the second iteration and adopted the CLEAN model of the iteration as the mean model image of the field of view for the observation period. The AD Leo emission at the phase centre was masked from the final model image and the model visibilities were subtracted from the corrected data. The remaining corrected visibilities pertain to purely stellar emission. This stellar data was subjected to two forms of analysis as described below.
@@
First, we developed a python based package, Visibility Averaged Dynamic spectrum (VISAD), to extract the visibilities from a calibrated CASA measurement set (MS) and generate a dynamic spectrum (DS) with required time and frequency averaging. The package has routines to perform additional data flagging on the calibrated MS, if required. The package uses CASA tasks inside a python environment and employs the parallel processing tools in python (multiprocessing111https://docs.python.org/3/library/multiprocessing.html#reference) enabling analysis of several hours of wide bandwidth datasets within a couple of minutes depending on the available machine cores. The module is made publicly available via GitHub222https://github.com/atul3790/Visibility-averaged-DS. We obtained both dynamic spectra and band averaged light curves in Stokes I and V using VISAD. The Stokes V data were corrected as per the suggestions in Das et al. (2020) to adhere to the standard IAU/IEEE convention.
@@
We have also produced image spectral cubes for the entire observing period with varying time and frequency averaging, carefully chosen to detect the star with at least a signal to noise ratio of 5 in every spectro-temporal slice. This exercise generated two data sets: a library of stellar spectra for different time bins spanning the entire observation period and a band averaged stellar light curve. The spectra were generated with a resolution of 20 MHz and a time averaging of 2 - 3 min during the very active periods. During periods of weak activity a 50 MHz averaging with 5 - 15 min time averaging was employed. We used the image cube information to cross-verify the flux densities obtained from visibility averaged DS for the flare period. The values matched within 15 - 20%. A source of this discrepancy is the additional flagging done in a couple of higher frequency channels while using VISAD. The presence of artefacts away from the source would also play a role in the observed difference in the flux density.
In this work, we will discuss the spectro-temporal nature of the observed activity, discuss the physical mechanism and infer physical properties of the emission region in the corona.
@@ 3.1 Results from VISAD
Figure 1a shows the Stokes V dynamic spectrum from VISAD made with 100 s and 10 MHz averaging in the top panel. The bottom panels shows the band averaged light curve obtained from VISAD in Stokes I and V, with 100 s averaging. Until 23:09:00 UT, we find multiple epochs of strong intermittent flaring activity, during which the stellar suddenly rises beyond 10 K with circular polarisation levels between 60 - 90%. Note that the values reported are averaged over the entire stellar disk. So the true will be higher depending on how small the area of the active region is compared to the size of the star. The flare rise phases are sudden and are followed by decay phases in both and polarisation fraction, as clearly seen in the band averaged light curves. We mark this initial 2 h period as a period of strong activity. Analysing the typical flux levels to which Stokes V emission falls back to after the sudden impulsive rise in the light curve, a basal Stokes V flux level is fixed at -2 mJy, corresponding to =-0.5 K. A similar exercise on the Stokes I light curve leads to a basal flux level of 4 mJy ( =1.0 K) which is about twice the typical quiescent flux seen during the quasi-steady period, as noted in Fig. 1. The basal flux levels are used to define flaring epochs during the strong activity period. A flaring epoch is said to begin when the stellar flux rises above the basal flux level and ends when the flux falls back below the basal level. Based on this definition, four flaring epochs (F1 - 4; see Fig. 1b-c) with flare decay timescales 5 - 30 min are identified in both of the Stokes light curves.
After the strong activity period, the star goes through 1.4 h period of weak activity between 2021-12-03 23:09:00 UT to 2021-12-04 00:24:00 UT (See, Fig. 1b-c). This period is characterised by high Stokes V 0.5 - 1.510, primarily the in higher frequency bands (700 MHz) as seen in VISAD spectrum (Fig. 1a). The observed Stokes V and Stokes I is much brighter than the expected coronal temperature of 3 MK based on X-ray studies (e.g., van den Besselaar et al., 2003; Namekata et al., 2020). Additionally, the band-averaged flux light curve in both Stokes I and V demonstrate high polarisation levels of 50 - 90% during the weak activity period, with a gradually varying profile resultant from the dynamic spectral structure of the emission. Such emission characteristics after flares have been observed in Sun as well and is interpreted as a signature of high-energy particles trapped in post-flare loops (Gergely, 1986; Gopalswamy, 2011; Hillaris et al., 2016; Morosan et al., 2019; Salas-Matamoros & Klein, 2020). We believe that the same phenomena is also behind the observed emission characteristics in this case as well.
After the weak activity period, we find that the Stokes V and I light curves tends to fluctuate around characteristic mean flux values of -0.5 and 1.5 mJy with relatively low circular polarisation levels. We define this phase as the quasi-steady phase.
@@ 3.2 Results from imaging analysis
To independently verify the flux density and polarisation trends presented in Section 3.1, we analysed the band-averaged light curve and image spectral cubes obtained from direct imaging (described in Section 2). Figure 2 shows the imaging-based light curve in Stokes I and V. Stokes V images have better SNR than those in Stokes I as evident from the lower error bars, due to relatively low levels of artefacts. Additionally, the number of galactic and extra-galactic sources with circular polarisation fraction exceeding are significantly low (e.g., Saikia & Salter, 1988; Macquart & Melrose, 2000; Enßlin et al., 2017; Callingham et al., 2023). Apart from the few time steps where only 5 upper limits could be determined either in V or I maps, we have good polarisation estimates across the light curve. As mentioned in Sect. 2, overall the fluxes from VISAD and imaging procedures agree within 20%. This flux uncertainty does not affect our analysis or physical modelling in Sect. 4. This is because, the emission being highly polarised in the flare and weak activity phase, is likely via some coherent emission processes. These coherent emission processes are often subject to a host of intrinsic physical processes, which are hard to nail down with just the available radio data (see, Bastian, 1990; Melrose, 2017b; Vedantham, 2021, for an overview). So our analysis in Sect. 4 requires absolute flux estimation accuracy only within an order of magnitude in (Bastian et al., 1990). The strong activity light curve matches well with the VISAD results and the significantly polarised weak activity period that followed is also well captured in the image-based light curve. During the weak activity period, Stokes I flux peaks at 2.5 mJy with a polarisation 60% similar to that seen in the VISAD light curve (Fig. 1(c)). A gradual decline phase after 00:40:00 UT is also seen in the Stokes I and V light curves (Fig. 2a-b), akin to the VISAD light curves (Fig. 1(b-c)).
In order to understand the spectral evolution of stellar emission, the stellar spectra derived from image spectral cubes described in Sect. 2 were analysed. Figure 3 shows example Stokes V spectra from specific time steps during the strong and weak activity period. Polarisation fraction is colour-coded. The spectrum shown in Fig. 3 for the strong activity period is chosen to coincide the peak epoch of the strongest flare (F3). The Stokes V flux of the star rises to around 16 mJy ( K) at high frequencies during this timestep (2021-02-03 22:12:13 UT) as seen in VISAD dynamic spectrum. Circular polarisation reaches to 100% during this time. The spectrum during the weak activity period shows an enhanced brightening with a significantly enhanced circular polarisation in the high frequency bands (700 MHz). This agrees with the observed enhanced confined to high frequency ranges in the VISAD dynamic spectrum during the weak activity phase (see, Fig. 1).

### 04 Mar 2024 v2

@@ Abstract
Context:AD Leo is a young and active M dwarf with high flaring rates across the X-ray-to-radio bands. Flares accelerate particles in the outer coronal layers and often impact exo-space weather. Wide-band radio dynamic spectra let us explore the evolution of particle acceleration activity across the corona. Identifying the emission features and modelling the mechanisms can provide insights into the possible physical scenarios driving the particle acceleration processes.
@@
Methods:The python-based module, Visibility Averaged Dynamic spectrum (VISAD), was developed to obtain the visibility-averaged wide-band dynamic spectra. Direct imaging was also performed with different frequency-time averaging. Based on existing observational results on AD Leo and on solar active region models, radial profiles of electron density and magnetic fields were derived. Applying these models, we explored the possible emission mechanisms and magnetic field profile of the flaring active region.
Results:The star displayed a high brightness temperature ( 10 – 10 K) throughout the observation. The emission was also nearly 100% left circularly polarised during bursts. The post-flare phase was characterised by a highly polarised (60 - 80%) solar-like type IV burst confined above 700 MHz.
Conclusions:The flare emission favours a Z -mode or a higher harmonic X-mode electron cyclotron maser emission mechanism. The 700 MHz post-flare activity is consistent with a type-IV radio burst from flare-accelerated particles trapped in magnetic loops, which could be a coronal mass ejection (CME) signature. This is the first solar-like type-IV burst reported on a young active M dwarf belonging to a different age - related activity population (‘C’ branch) compared to the Sun (‘I’ branch). We also find that a multipole expansion model of the active region magnetic field better accounts for the observed radio emission than a solar-like active region profile.
@@
Low-mass main-sequence stars of spectral type M (M dwarfs) make up the most numerous stellar type in our galaxy. They also have the highest occurrence rate of Earth-like exoplanets in their habitable zones (Bashi et al., 2020). Meanwhile, M-dwarfs form the most active population among cool main-sequence stars with relatively high values of flaring rates, X-ray-to-bolometric flux ratio, and other activity metrics (see Donati & Landstreet, 2009 for a review). Moreover, the younger stars are more active with stronger magnetic fields and faster rotation rates (e.g. Vidotto et al., 2014; Lund et al., 2020). Stellar flares cause energetic space-weather events including the release of energetic particles, intense high-energy radiation and large-scale eruptions, which threaten the stability of planetary atmospheres and hence the habitability. So, it is important to understand the mechanism, evolution, and space-weather consequences of M dwarf flares.
@@
Flare emission in the metrewave band is sensitive to various coherent emission signatures from flare-accelerated electrons either trapped in low-coronal magnetic field structures or streaming across coronal layers (e.g. Wild, 1970; McLean & Labrum, 1985; Bastian, 1990; Melrose, 2009, 2017a; Zic et al., 2020; Vedantham, 2021). The two prominent coherent emission mechanisms discussed in the literature include the electron cyclotron maser emission (ECME) and the plasma emission (Dulk, 1985; Güdel, 2002; Melrose, 2017a). ECME dominates when the ratio of plasma frequency to gyro-frequency (/) at the flare source region is less than 1.5 (Melrose & Dulk, 1982; Melrose et al., 1984a; Dulk, 1985; Dulk & Winglee, 1987). When /3, the favoured emission mechanism is plasma emission (e.g. Ginzburg & Zhelezniakov, 1958; Tsytovich & Kaplan, 1969; Melrose & Sy, 1972; Cairns & Melrose, 1985; Melrose, 2009). Depending on the emission mechanism, the coherent emission frequencies correspond to either the fundamental mode or the harmonics of the local plasma frequency () or gyro-frequency (), letting us infer either the local density or magnetic field strength, respectively. Apart from the flare emission, M dwarfs are also known to display non-flaring (or non-impulsive) continuum emission with a high brightness temperature () (e.g. White et al., 1989; Bastian, 1990). In a few GHz (¡2 GHz) bands, the continuum ranges from 10 - 10 K (Osten et al., 2008; Villadsen & Hallinan, 2019). In metrewavebands, can be even higher (e.g. Bastian, 1994; Güdel, 2002; Villadsen & Hallinan, 2019). The continuum emission can form either from a gyrosynchrotron mechanism, or a combination of various coherent emission processes operating at multiple active regions (Bastian, 1996; Benz, 1996; Zaitsev et al., 1997).
@@
There have been many studies that explored radio emission from AD Leo, reporting several flaring events. While instrumental limitations forced earlier studies to use only a few discontinuous narrow frequency bands (e.g. Spangler et al., 1974; Bastian et al., 1990), recent observations with continuous wide-band instruments have enabled spectroscopic explorations revealing rich variability in the dynamic spectrum (e.g. Osten et al., 2008; Villadsen & Hallinan, 2019; Zhang et al., 2023). These studies have revealed several features such as fast pulsations at scales of ms - min and high drift rates in the dynamic spectrum, often related to the propagation of energetic particles or spatio-temporal evolution of plasma instabilities in the associated active regions (e.g. Bastian et al., 1990; Osten & Bastian, 2006; Osten et al., 2008; Zhang et al., 2023). However, much of these observations had been close to or above 1 GHz. Here, we present an 8 h monitoring of AD Leo across a wide sub-Giga Hertz band, which is expected to probe higher atmospheric layers than those probed by Giga Hertz bands. Though the original study was planned with simultaneous observation spanning from the X-ray to the radio band, only the radio observations were successful due to various technical issues. We present the results from metrewave observations here.
@@ 2 Data and methodology
AD Leo was observed with uGMRT (Gupta et al., 2017) in Band 4 (550 - 850 MHz) from 3 December 2021, 21:17:41 UT to 4 December 2021, 04:17:41 UT at 5 s and 100 kHz time and frequency resolutions, respectively. The data were gathered in RR and LL polarisations, letting us explore the spectro-temporal evolution of the circular polarisation along with brightness temperature () or flux density. The data analysis was done using the functionalities in the interferometric analysis package, Common Astronomy Software Applications (CASA; CASA Team et al., 2022). The data were first flagged and calibrated using standard procedures. The flux calibrator was 3C147, and the phase calibrator was 1021+219. Portions of data affected by radio frequency interference from terrestrial sources and systematic instrumental effects were identified and flagged manually. The automatic flagging routine in CASA, namely tfcrop, was also used. This was followed by bandpass calibration using the flux calibrator, and then by time-dependent gain calibration using the phase calibrator. The bandpass solution and the time-dependent gain solution were then extrapolated to the target star field. The calibrated target field data were further flagged using the Rflag routine in CASA. Initial imaging was done using the CASA task tclean in mtmfs mode, with nterms set to two and averaging over the entire observing time. This produced a mean model image for the entire spectral band, which was taken as the initial model, and self-calibration was performed in an iterative manner (Cornwell & Fomalont, 1999) to improve it. After the second iteration, an optimal model was obtained. So, we stopped at the second iteration and adopted the CLEAN model of the iteration as the mean model image of the field of view for the observation period. The AD Leo emission at the phase centre was masked from the final model image, and the model visibilities were subtracted from the corrected data. The remaining corrected visibilities pertain to purely stellar emission. This stellar data were subjected to two forms of analysis as described below.
@@
First, we developed a python-based package, Visibility Averaged Dynamic spectrum (VISAD), to extract the visibilities from a calibrated CASA measurement set (MS) and generate a dynamic spectrum (DS) with required time and frequency averaging. The package has routines to perform additional data flagging on the calibrated MS if required. The package uses CASA tasks inside a python environment and employs the parallel processing tools in python (multiprocessing111https://docs.python.org/3/library/multiprocessing.html#reference) enabling an analysis of several hours of wide-bandwidth datasets within a couple of minutes depending on the available machine cores. The module is made publicly available via GitHub222https://github.com/atul3790/Visibility-averaged-DS. We obtained both dynamic spectra and band-averaged light curves in Stokes I and V using VISAD. The Stokes V data were corrected as per the suggestions in Das et al. (2020) to adhere to the standard IAU/IEEE convention.
@@
We also produced image spectral cubes for the entire observing period with varying time and frequency averaging, which were carefully chosen to detect the star with a signal-to-noise ratio (S/N) of at least five in every spectro-temporal slice. This exercise generated two data sets: a library of stellar spectra for different time bins spanning the entire observation period and a band-averaged stellar light curve. The spectra were generated with a resolution of 20 MHz and a time averaging of 2 - 3 min during the very active periods. During periods of weak activity, a 50 MHz averaging with 5 - 15 min time averaging was employed. We used the image cube information to cross-verify the flux densities obtained from a visibility-averaged DS for the flare period. The values matched within 15 - 20%. A source of this discrepancy is the additional flagging done in a couple of higher frequency channels while using VISAD. The presence of artefacts away from the source would also play a role in the observed difference in the flux density. In this work, we discuss the spectro-temporal nature of the observed activity, the possible physical mechanism, and infer physical properties of the emission region in the corona.
@@ 3.1 Results from VISAD
Figure 1a shows the Stokes V dynamic spectrum from VISAD made with 100 s and 10 MHz averaging in the top panel. The bottom panels shows the band-averaged light curve obtained from VISAD in Stokes I and V, with 100 s averaging. Until 23:09:00 UT, we find multiple epochs of strong intermittent flaring activity, during which the stellar suddenly rises beyond 10 K with circular polarisation levels between 60 and 90%. We note that the values reported are averaged over the entire stellar disc, thus the true will be higher depending on how small the area of the active region is compared to the size of the star. The flare rise phases are sudden and are followed by decay phases in both and polarisation fraction, as clearly seen in the band-averaged light curves. We mark this initial 2 h period as a period of strong activity. Analysing the typical flux levels to which Stokes V emission falls back after the sudden impulsive rise in the light curve, a basal Stokes V flux level is fixed at -2 mJy, corresponding to =-0.5 K. A similar exercise on the Stokes I light curve leads to a basal flux level of 4 mJy ( =1.0 K), which is about twice the typical quiescent flux seen during the quasi-steady period, as noted in Fig. 1. The basal flux levels are used to define flaring epochs during the strong activity period. A flaring epoch is said to begin when the stellar flux rises above the basal flux level and ends when the flux falls back below the basal level. Based on this definition, four flaring epochs (F1 - 4; see Fig. 1b-c) with flare decay timescales of 5 - 30 min are identified in both of the Stokes light curves.
After the strong activity period, the star goes through 1.4 h period of weak activity between 2021-12-03 23:09:00 UT and 2021-12-04 00:24:00 UT (see Figs. 1b-c). This period is characterised by high Stokes V 0.5 - 1.510, primarily the in higher frequency bands (700 MHz) as seen in the VISAD spectrum (Fig. 1a). The observed Stokes V and Stokes I is much brighter than the expected coronal temperature of 3 MK based on X-ray studies (e.g. van den Besselaar et al., 2003; Namekata et al., 2020). Additionally, the band-averaged flux light curves in both Stokes I and V demonstrate high polarisation levels of 50 - 90% during the weak activity period, with a gradually varying profile resultant from the dynamic spectral structure of the emission. Such emission characteristics after flares have been observed in the Sun as well and are interpreted as signatures of high-energy particles trapped in post-flare loops (Gergely, 1986; Gopalswamy, 2011; Hillaris et al., 2016; Morosan et al., 2019; Salas-Matamoros & Klein, 2020). We believe that the same phenomenon is also behind the observed emission characteristics in this case.
After the weak activity period, we find that the Stokes V and I light curves tend to fluctuate around characteristic mean flux values of -0.5 and 1.5 mJy with relatively low circular polarisation levels. We define this phase as the quasi-steady phase.
@@ 3.2 Results from imaging analysis
To independently verify the flux density and polarisation trends presented in Section 3.1, we analysed the band-averaged light curve and image spectral cubes obtained from direct imaging (described in Section 2). Figure 2 shows the imaging-based light curve in Stokes I and V. Stokes V images have better S/N than those in Stokes I as evident from the lower error bars, due to relatively low levels of artefacts. Additionally, the number of galactic and extra-galactic sources with a circular polarisation fraction exceeding are significantly low (e.g. Saikia & Salter, 1988; Macquart & Melrose, 2000; Enßlin et al., 2017; Callingham et al., 2023). Apart from the few time steps where only 5 upper limits could be determined either in V or I maps, we have good polarisation estimates across the light curve. As mentioned in Sect. 2, overall the fluxes from VISAD and imaging procedures agree within 20%. This flux uncertainty does not affect our analysis or physical modelling in Sect. 4. This is because, the emission being highly polarised in the flare and weak activity phase, is likely via some coherent emission processes. These coherent emission processes are often subject to a host of intrinsic physical processes, which are hard to nail down with just the available radio data (see, Bastian, 1990; Melrose, 2017b and Vedantham, 2021 for an overview). Our analysis in Sect. 4 thus requires absolute flux estimation accuracy only within an order of magnitude in (Bastian et al., 1990). The strong activity light curve matches well with the VISAD results, and the significantly polarised weak activity period that followed is also well captured in the image-based light curve. During the weak activity period, Stokes I flux peaks at 2.5 mJy with a polarisation of 60% similar to that seen in the VISAD light curve (Fig. 1(c)). A gradual decline phase after 00:40:00 UT is also seen in the Stokes I and V light curves (Fig. 2a-b), akin to the VISAD light curves (Fig. 1(b-c)).
In order to understand the spectral evolution of stellar emission, the stellar spectra derived from image spectral cubes described in Sect. 2 were analysed. Figure 3 shows example Stokes V spectra from specific time steps during the strong and weak activity period. Polarisation fraction is colour-coded. The spectrum shown in Fig. 3 for the strong activity period is chosen to coincide with the peak epoch of the strongest flare (F3). The Stokes V flux of the star rises to around 16 mJy ( K) at high frequencies during this timestep (2021-02-03 22:12:13 UT) as seen in VISAD dynamic spectrum. Circular polarisation reaches 100% during this time. The spectrum during the weak activity period shows an enhanced brightening with a significantly enhanced circular polarisation in the high-frequency bands (700 MHz). This agrees with the observed enhanced confined to high-frequency ranges in the VISAD dynamic spectrum during the weak activity phase (see Fig. 1).